You are trying to decide between putting $3000 or $4000 annually for the next 20 years into an investment yielding 7%. What is the difference in the value of investing the extra $1000 for 20 years (ordinary annuity)? What is the future value if the payment is deposited at the beginning of each year (annuity due)?
Answer #1

The difference in the value of investing the extra $1000 for 20 years ordinary annuity at 7% = 4000*((1+7%)20-1)/7% -3000 (1+7%)20-1)/7% = 40,995.49

The FV if annuity due = (1+7%) * 1000* (1+7%)20-1)/7% = 43,865.18

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well